OpenStudy (faman39):

Please help with my math prob! i hav a exams, tats y i need to understand those prob and hav to know how those prob can solve easily before exams coz if i dnt, i wnt able to pass. So pls help with all. I will give medals for all to them, who will able to explain wit correct answer Rebekah purchased a sweater that was reduced by 25 % (or 1/4 ) of its original price. She also used a store coupon for $5 . When she made the purchase, she paid the cashier $19 . What was the original price of the sweater?

hero (hero):

This is a problem where you have to work backward. Rebekah paid the cashier 19 dollars. Without the $5 coupon, it would have cost her 19 + 5 = 24 dollars. The 24 dollars is the discount price. The general formula for finding the discount price of an item is x - %x = discount price, and x = the original price. If we input the given values % = 1/4 or.25, and 24 = discount price, then we have: x - .25x = 24 Solving for x, we get x(1-.25) = 24 x(.75) = 24 x = 24/.75 x = 32 Now remember, x equal the original price, therefore, 32 dollars is the original price of the sweater.